Fawn lives on a hill. She goes to work by bicycle at a speed of 20 mph. Work is 10 miles away. She returns home at a speed of 15 mph along the same 10 mile route. How long does she spend commuting every day?
Answer by checkley77(12844)

D=RT
10=20T
T=10/20
T=.5 HOURS OR 30 MINUTES.
10=15T
T=10/15
T=2/3 HOURS OR 40 MINUTES.
30+40=70 MINUTES OR 1 HR. & 10 MINUTES FOR THE ROUND TRIP COMMUTE.